OpenStudy (lukecrayonz):

(x-1/2x-4)-(x+2/3x)=1 (x-1/2x-4)-(x+2/3x)=1 @Mathematics

OpenStudy (anonymous):

(x-1/2x-4)-(x+2/3x)=1 =>Possible intermediate steps: -7/(6 x)-4 = 1 Add 4 to both sides: -7/(6 x) = 5 Divide both sides by -7/6: 1/x = -30/7 Raise both sides to the power of -1: x = -7/30

OpenStudy (lukecrayonz):

Thank you :D Please help me with the other one, since I honestly have no idea where to start D: